OpenStudy (anonymous):

Last question...Factor completely. x^4-50x^2y^2+49y^4

OpenStudy (ranga):

x^4-50x^2y^2+49y^4 = x^4 - x^2y^2 - 49x^2y^2 + 49y^4 Factor the first two and the last two and write them as a product of two factors. Then you can factor each of them again with the identity: a^2 - b^2 = (a + b)(a - b)

OpenStudy (anonymous):

Ok...what I started with was (x^2+y^2)(x^2-49y^2) (x^2+y^2)(x-7y)(x+7y) Did I go wrong somewhere?

OpenStudy (ranga):

First line should be: (x^2- y^2)(x^2 - 49y^2)